On November 1st, you will buy a Lamborghini for $300,000. You will pay $150,000 in cash at the time of purchase and borrow the remaining amount from the bank. This loan will be paid in five equal annual installments that include both the principle and 8 percent interest on the declining balance.

a) determine the amount of the annual payment

b) determine the total dollars of interest that you will pay for this loan

C) Determine the amount of interest that is included in the first payment.
Answer #1

Question - (a)

Loan amount = 300,000 - 150,000 = 150,000

Annual payment = Loan amount / Present value of annuity

PVA = [ 1 - (1+r)-n ] / r

r = 0.08

n = 5 years

= [ 1 - (1.08)-5 ] / 0.08

= 3.9927100371

Annual payment = 150,000 / 3.9927100371 = 37568.47

Question - (b)

Total dollars of interest paid = Total Installments paid - Principal loan amount

= 5 * 37568.47 - 150,000

= 37,842.35

Note : What ever we paid in excess of Principal loan amount is for Interest.

Question - (c)

Interest included in first payment = 150,000 * 8% = 12000